. (5 pts) A sample of 27.0 mL of 0.100 M HBr reacts with 45.3 mL. of Ba(OH)2 Find the concentration (molarity) of Ba(OH)2. Please use the gfu approach and show all work for full credit. If you just show a numerical answer, you will receive only 0.25 pt. 2 HBr + Ba(OH)--> BaBr2 + 2 H2O
Answer #1

The reaction is 2HBr+ Ba(OH)2 ------->BaBr2+2H2O

2 moles of HBr requires 1 mole of Ba(OH)2

moles= molarity* Volume in liters, 1000ml= 1L

moles of HBr= 0.1*27/1000=0.0027

hence moles of Ba(OH)2 required= moles of HBr/2= 0.0027/2=0.00135

volume of Ba(OH)2= 45.3 ml, converting this into liters, volume of Ba(OH)2= 45.3/1000=0.0453L

Hence concentration of Ba(OH)2= moles/Volume in liters= 0.00135/0.0453 moles/L=0.0298M